9) It is known that the distance between nucleotides in DNA strands is 34 × 10-11m. How long is the 287 amino acid hemoglobin gene?


1
Expert's answer
2021-02-18T12:50:10-0500

Ans. Distance between nucleotides in DNA strands is 34 × 10-11m


We know that 1 amino acid is made up of 1 codon.

1 codon is made up of 3 nucleotide bases.

Thus, if we have 287 amino acids then, number of nucleotides will be: 287 *3= 861

Now since its already mention that distance between 2 nucleotides in DNA strand is 34 * 10-11m

Hence the distance or length of 287 amino acid hemoglobin gene or we can also say as the length of 861 nucleotides containing gene is:

(861-1) * 34 * 10-11 = 29240 * 10-11m


  • Note: here we did 861-1 since we are calculating distance between nucleotides, 34*10-11m is distance between 2 nucleotides, so we need to do n-1 here, where n is number of nucleotides.


Thus, length of 287 amino acid hemoglobin gene is 29240 * 10-11m.